OpenStudy (anonymous):

find the LCM of 20 and 30... what is the LCM anyways?

OpenStudy (anonymous):

least common multiple. 60 is the LCM of 20 and 30, because it is the smallest number that is an even multiple of both of them.

OpenStudy (anonymous):

thank you

OpenStudy (ybarrap):

LCM = (num1 * num2) / GCD (num1,num2) But this doesn't tell you what LCM means Take the unique prime factors of num1 and num2 together multiply them. This formula just gives a way to get that number directly For example, 20*30/GCD(20,30) = 20*30/10 = 2*30 = 60. Factors of 20 = 2^2*5 Factors of 30 = 2*3*5 Factors in common (GCD): 2,5 Other factors: 2,3 All unique factors: 2,2,3,5. Their product is 2*2*3*5 = 60